# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v1

@@ B.4 Training Details
We perform closed-loop fine-tuning on selected modules of the IL pre-trained trajectory generation model (Pluto). As shown in the ablation results (Section 5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on 2×2$\times$2 × Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table 6 and Table 6.
@@ B.5 Baselines Detailed Description
To comprehensively evaluate RIFT in an AV-centered closed-loop simulation environment, we compare it against a range of baselines, including pure imitation learning (IL), pure reinforcement learning (RL), and various fine-tuning approaches based on IL, RL, or their combination. We initialize all fine-tuning methods from the pre-trained Pluto checkpoint and fine-tune only the trajectory scoring head to preserve trajectory-level realism. The details of each baseline are summarized below.
@@
  • Pluto [53] is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In AV-centered closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
@@
  • FREA [14] is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centered closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
@@
  • PPO [14] is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
@@
  • FPPO-RS [14] is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
@@
  • PPO-Pluto fine-tunes the pre-trained trajectory generation model using the PPO algorithm [58]. The fine-tuning follows the same reward structure as detailed in Section B.6, aligning with RIFT.
@@
  • REINFORCE-Pluto employs the REINFORCE algorithm [59] to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section B.6.
@@
  • GRPO-Pluto utilizes the basic GRPO algorithm [20] for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
@@
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite [69] serves as the expert model, providing supervision at the target speed level.
@@
  • RTR-Pluto [21] is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
@@
  • RS-Pluto [22] also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
B.6 State-Wise Reward Model Setup
@@
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following [73]. Different styles are constructed by combining weights assigned to each reward component (detailed in Table 6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
@@
Building on the reward definitions above, we construct a state-wise reward model $\mathrm{StateWiseRM}(\cdot)$, which computes a scalar reward based on a set of interpretable features extracted from each trajectory point $\tau^{t}_{i}$. Specifically, we define a feature extraction function $\phi(\tau^{t}_{i})$ as:
$\phi(\tau^{t}_{i})=(\mathbbm{1}_{\text{collision}},\mathbbm{1}_{\text{boundary}},a_{\text{long}},a_{\text{lat}},\theta_{f},x_{f},v,a),$ (21)
@@
  • $\theta_{f}$ is the heading deviation with respect to the lane direction.
@@
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\tau^{t}_{i})).$ (22)
All features, except the infraction indicators, are directly derived from the trajectory. To estimate future infractions, we follow the forecasting model in [69] to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected using the Euclidean Signed Distance Field (ESDF) [53].
@@
B.7 Controllability and Realism Metrics
Controllability. Following the metric design principles proposed in the WOSAC challenge [34] and other widely adopted evaluation frameworks [42, 63], we adopt a set of well-established metrics to comprehensively assess different aspects of controllability:
  • Scenario Collision Per Kilometer (CPK) [63]: the average number of scenario collisions per kilometer of driving distance.
  • Off-Road Rate (ORR) [63]: the percentage of time that CBVs spend off-road on average.
  • Route Progress (RP) [63]: the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) [64]: the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the immediate interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) [65] : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
Realism. Following [63], most realism metrics rely on ground-truth trajectories of human drivers to evaluate how closely simulation agents mimic real-world behavior. However, since all our experiments are conducted in the CARLA simulator without expert demonstrations, we adopt distribution-level metrics to assess the realism of CBV behavior in terms of speed and acceleration. Specifically, we employ the following two metrics:
  • Wasserstein Distance (WD) [68]: measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
@@ B.7 Controllability and Realism Metrics
  • Shapiro–Wilk test (SW) [66]: assesses whether the distribution of speed and acceleration follows a normal distribution, capturing the statistical naturalness of CBV motion.
@@ C.1 AV Methods Implementation
To assess the effectiveness of RIFT in generating reliable and interactive scenarios for AV evaluation in the AV-centered closed-loop simulation environment, we evaluate the following representative and stable AV methods:
  • PDM-Lite [69]: A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT [71]: An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
@@ C.2 AV Evaluation Metrics
As detailed in Sections B.1 and B.4, we develop an AV-centered closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section B.1, replacing predefined scenarios with CBV-generated traffic flows precludes the evaluation of specific AV capabilities. Since our primary objective is to validate the effectiveness of RIFT in providing reliable traffic scenarios for AV evaluation, we follow the default evaluation metrics from CARLA Leaderboard to assess the overall driving performance of the AV. Additionally, we introduce a new metric to quantify the AV blockage events, which may arise when traffic flows are unreliable. The evaluation metrics are summarized as follows:
@@
  • Route Completion (RC): $R_{i}$ — The percentage of the route distance completed by the agent. The maximum value is 100.
  • Infraction penalty (IP): $P_{i}=\prod_{j}p_{j}^{\text{infractions}_{j}}$ — This metric tracks various types of infractions and aggregates all instances triggered by the agent as a geometric series. The agent starts with an ideal base score of 1.0, which decreases with each infraction to a minimum of 0.
@@ Appendix D Additional Results
D.1 Detailed Qualitative Results of Controllability
As discussed in Section 5.4, we investigate the controllability of RIFT under different reward configurations. The aggressive variant applies a reduced collision penalty and places greater emphasis on driving efficiency (Table 6), encouraging assertive behaviors such as overtaking. In contrast, the normal configuration imposes a higher collision penalty to promote safer and more conservative driving behaviors.
Quantitative results in Table 3 show that the aggressive variant achieves higher Progress and Speed Realism scores, albeit at the cost of increased infractions. To complement these findings, Figure 6 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving behaviors according to user-specified reward configurations.
Figure 6: Qualitative illustration of RIFT’s controllability under different reward configurations. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
@@ D.2 Detailed Analysis in Driving Comfort
  • Driving Jerk (Jerk): the time derivative of acceleration, quantifying the abruptness of acceleration changes and the smoothness of CBV trajectories.
@@
$\begin{split}\text{Frame Variable Smoothness (FVS)}&=\begin{cases}\text{True}&\text{if lower bound}\leq p_{\,i}\leq\text{upper bound}.\\ \text{False}&\text{otherwise}\end{cases}\\ &p\in\text{smoothness vars},\,0\leq i\leq\text{total frames}\end{split}$ (25)
@@
Method PDM-Lite [69] PlanT [71]
@@
Pluto [53] 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO [14] 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA [14] 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS [14] 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
@@
RS-Pluto [22] 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto [21] 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
@@
Main Results. The quantitative results of the comfort metrics are presented in Table 7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section 5.2 achieve relatively lower discomfort levels, a high baseline of discomfort persists across methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure 7, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise—predicting only the immediate next action—the reference speed may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment—RIFT adopts a different approach. It employs a state-wise reward model (see Section B.6) that quantifies comfort within the trajectory’s virtual time domain.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure 7). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table 7.
@@
D.3 Visualization of the AV-Centered Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centered scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure 8, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
The visualizations demonstrate the ability of RIFT to generate temporally coherent, realistic, and controllable trajectories across a variety of traffic situations. Even under complex and evolving closed-loop conditions, RIFT maintains stable multimodal behavior, highlighting its effectiveness in simulating realistic and controllable traffic flows around the AV.
Figure 8: Extra temporal visualization of RIFT in the AV-centered closed-loop simulation environment. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.

### v2

@@ B.4 Training Details
We perform closed-loop fine-tuning on selected modules of the IL pre-trained planning model (Pluto). As shown in the ablation results (Section˜5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on $2\times$ Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table˜6 and Table˜6.
@@ B.5 Baselines Detailed Description
To comprehensively evaluate RIFT in an AV-centric closed-loop simulation environment, we compare it against a range of baselines, including pure imitation learning (IL), pure reinforcement learning (RL), and various fine-tuning approaches based on IL, RL, or their combination. We initialize all fine-tuning methods from the pre-trained Pluto checkpoint and fine-tune only the trajectory scoring head to preserve trajectory-level realism. The details of each baseline are summarized below.
@@
  • Pluto cheng2024pluto is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In the AV-centric closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
@@
  • FREA chen2024frea is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centric closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
@@
  • PPO chen2024frea is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
@@
  • FPPO-RS chen2024frea is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
@@
  • PPO-Pluto fine-tunes the pre-trained planning model using the PPO algorithm PPO . The fine-tuning follows the same reward structure as detailed in Section˜B.7, aligning with RIFT.
@@
  • REINFORCE-Pluto employs the REINFORCE algorithm sutton1999reinforce to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section˜B.7.
@@
  • GRPO-Pluto utilizes the basic GRPO algorithm shao2024grpo for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
@@
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite Beibwenger2024pdmLite serves as the expert model, providing supervision at the target speed level.
@@
  • RTR-Pluto zhang2023rtr is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
@@
  • RS-Pluto peng2024improving also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
B.6 Forward Simulation
Trajectory-based imitation learning often neglects underlying system dynamics, leading to discrepancies between planned and executed behavior cheng2024planTF . To address this, we perform a closed-loop forward simulation for each candidate trajectory ${\tau}_{i}$, yielding a rollout $\widetilde{\tau}_{i}$. The simulation integrates a PID controller for trajectory tracking with a kinematic bicycle model for state propagation. The PID controller is identical to that used during closed-loop execution, ensuring behavioral consistency between training and deployment, while the simulation time step is aligned with the temporal resolution of candidate trajectories. By evaluating rollouts rather than raw candidate trajectories, we mitigate this gap and obtain more reliable assessments.
B.7 State-Wise Reward Model Setup
@@
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following cusumano2025robust . Different styles are constructed by combining weights assigned to each reward component (detailed in Table˜6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
@@
Building on the reward definitions above, we construct a state-wise reward model $\mathrm{StateWiseRM}(\cdot)$, which computes a scalar reward based on a set of interpretable features extracted from each rollout point $\widetilde{\tau}^{t}_{i}$. Specifically, we define a feature extraction function $\phi(\widetilde{\tau}^{t}_{i})$ as:
$\phi(\widetilde{\tau}^{t}_{i})=(\mathbbm{1}_{\text{collision}},\mathbbm{1}_{\text{boundary}},a_{\text{long}},a_{\text{lat}},\theta_{f},x_{f},v,a),$ (21)
@@
  • $\theta_{f}$ is the heading deviation concerning the lane direction.
@@
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\widetilde{\tau}^{t}_{i})).$ (22)
All features, except the infraction indicators, are directly derived from the rollouts. To estimate future infractions, we follow the forecasting model in Beibwenger2024pdmLite to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected using the Euclidean Signed Distance Field (ESDF) cheng2024pluto .
@@
B.8 Controllability and Realism Metrics
Kinematic Metrics. Following montali2023wosac , kinematic realism is typically evaluated against ground-truth trajectories. As CARLA provides no expert demonstrations, we adopt distribution-level metrics chen2024review ; huang2025IntSim to assess CBV behavior in terms of speed and acceleration. Specifically, we employ three measures—the Shapiro–Wilk test on speed (S-SW), the Wasserstein Distance on speed (S-WD), and the Shapiro–Wilk test on acceleration (A-SW)—defined as follows:
  • Wasserstein Distance (WD) vaserstein1969wd : measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
@@ B.8 Controllability and Realism Metrics
  • Shapiro–Wilk test (SW)shapiro1965sw : evaluates the normality of speed and acceleration distributions—a simplifying assumption supported by empirical traffic studieszhang2023rtr ; yan2023NeuralNDE —to capture the statistical naturalness of CBV motion.
@@
Interaction Metrics. Following the metric design principles proposed in the WOSAC challenge montali2023wosac and other widely adopted evaluation frameworks lin2024ccdiff ; chen2024review , we adopt a set of well-established metrics to comprehensively evaluate agent interactions:
  • Collision Per Kilometer (CPK) chen2024review : the average number of scenario collisions per kilometer of driving distance.
  • Route Progress (RP) chen2024review : the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) guo20232dttc : the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the immediate interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) venthuruthiyil2022act : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
Map Metrics. Map metrics evaluate adherence to road geometry, reflecting how well agents remain within drivable areas and comply with map constraints.
  • Off-Road Rate (ORR) chen2024review : the percentage of time that CBVs spend off-road on average.
@@ C.1 AV Methods Implementation
To assess the effectiveness of RIFT in generating reliable and interactive scenarios for AV evaluation in the AV-centric closed-loop simulation environment, we evaluate the following representative and stable AV methods:
  • PDM-Lite Beibwenger2024pdmLite : A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT Renz2022PlanT : An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
  • UniAD hu2023_uniad : A planning-oriented unified framework integrating perception, prediction, mapping, and planning into one end-to-end model using query-based interfaces.
  • VAD jiang2023vad : A fast, end-to-end vectorized driving paradigm representing scenes with vectorized motion and map elements for efficient, safe planning.
@@ C.2 AV Evaluation Metrics
As detailed in Sections˜B.1 and B.4, we develop an AV-centric closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section˜B.1, replacing predefined scenarios with CBV-generated traffic scenarios precludes the evaluation of specific AV capabilities. To systematically assess the quality of traffic scenarios generated by different CBV methods, we follow the practice of KING king and introduce PDM-Lite Beibwenger2024pdmLite —a rule-based privileged planner—as a reference AV. By measuring its performance under various CBV methods, we evaluate:
  • Feasibility, via PDM-Lite’s Driving Score (DS)—a high DS indicates the PDM-Lite can complete its route without severe collisions or rule violations, implying the generated traffic scenario is feasible.
  • Naturalness, via a newly proposed metric, Blocked Rate (BR)—a low BR suggests that CBVs do not unrealistically obstruct the AV, reflecting naturalistic behavior.
These metrics enable a principled comparison of traffic quality generated by different CBV methods. Furthermore, to assess the capacity of generated traffic scenarios to expose AV limitations, we test multiple learning-based AV methods under an identical CBV method and quantify their relative performance drop compared to PDM-Lite Beibwenger2024pdmLite . The relative driving score degradation ($\Delta{DS}$) reflects how effectively the traffic scenario stresses the AV policy, with larger drops indicating stronger capability in revealing planning weaknesses.
The evaluation metrics are summarized as follows:
@@
  • Relative Driving Score Degradation ($\Delta{DS}$): The reduction in Driving Score of a learning-based AV compared to PDM-Lite under the same CBV method, indicating how effectively the scenario reveals weaknesses in AV planning.
@@ Appendix D Additional Results
D.1 Detailed Qualitative Results of Style-Level Controllability
As discussed in Section˜5.4, we investigate the style-level controllability of RIFT under different reward configurations. The aggressive variant applies a reduced collision penalty and places greater emphasis on driving efficiency (Table˜6), encouraging assertive behaviors such as overtaking. In contrast, the normal configuration imposes a higher collision penalty to promote safer and more conservative driving behaviors.
Quantitative results in Table˜3 show that the aggressive variant achieves greater driving efficiency at the expense of more frequent collisions and off-road events. To complement these findings, Figure˜5 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving style according to user-specified reward configuration.
Figure 5: Qualitative illustration of RIFT’s style-level controllability under different reward configurations. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
@@ D.2 Detailed Analysis in Driving Comfort
  • Driving Jerk (Jerk): the time derivative of acceleration, quantifying the abruptness of acceleration changes and the smoothness of CBV rollouts.
@@
[…]
@@
Method PDM-Lite Beibwenger2024pdmLite PlanT Renz2022PlanT
@@
Pluto cheng2024pluto 56.45 $\pm$ 4.14 -0.16 $\pm$ 3.72 50.26 $\pm$ 2.17 -0.42 $\pm$ 3.38
PPO chen2024frea 74.76 $\pm$ 2.71 -0.51 $\pm$ 4.61 74.90 $\pm$ 1.21 0.40 $\pm$ 4.83
FREA chen2024frea 72.40 $\pm$ 1.72 0.29 $\pm$ 4.61 73.48 $\pm$ 3.83 -0.15 $\pm$ 4.91
FPPO-RS chen2024frea 68.33 $\pm$ 1.90 -0.07 $\pm$ 3.96 66.67 $\pm$ 0.82 -0.15 $\pm$ 3.95
@@
RS-Pluto peng2024improving 70.31 $\pm$ 4.07 0.32 $\pm$ 4.12 65.18 $\pm$ 2.11 -0.16 $\pm$ 4.07
RTR-Pluto zhang2023rtr 55.58 $\pm$ 4.76 -0.19 $\pm$ 3.37 45.12 $\pm$ 2.66 -0.14 $\pm$ 3.34
@@
Main Results. The quantitative results of the comfort metrics are presented in Table˜7. All CBV methods exhibit notable levels of driving discomfort. Although the more conservative methods identified in Section˜5.2 achieve relatively lower discomfort levels, a high baseline of discomfort persists across methods.
To investigate the underlying causes of discomfort, we further decouple the planned trajectories from the executed control actions. In CARLA, most CBV methods rely on PID controllers to transform high-level trajectory waypoints into executable driving commands, including throttle, steering, and brake. As shown in Figure˜6, the upper panel illustrates the speed tracking curve, while the middle panel presents the raw throttle signal and corresponding acceleration profile.
Because trajectory generation is performed state-wise—predicting only the immediate next action—the reference states may vary discontinuously over time. These discontinuities are amplified by the PID controller, whose binary throttle/brake responses induce abrupt changes in acceleration, ultimately leading to discomfort during vehicle operation. Such execution-level instabilities are a major contributor to the discomfort observed across CBV methods.
While many CBV methods attempt to mitigate discomfort through fine-tuning strategies that incorporate post-action feedback via reward shaping or expert action alignment—RIFT adopts a different approach. It employs a state-wise reward model (see Section˜B.7) that quantifies comfort within the trajectory’s virtual forward simulation.
To further analyze this, we visualize both the actual acceleration after executing a selected trajectory and the corresponding virtual-time acceleration (shown in Figure˜6). The results reveal that while virtual-time acceleration aligns with actual motion at the beginning of the trajectory, it underestimates acceleration variations in later segments. This leads to an overly conservative estimation of trajectory-level discomfort, resulting in insufficient supervision during training and reflected in RIFT’s comfort performance in Table˜7.
@@
D.3 Visualization of the AV-Centric Closed-Loop Simulation
To qualitatively evaluate the robustness of RIFT across diverse AV-centric scenarios, we provide additional temporal visualizations of closed-loop simulations. As shown in Figure˜7, the traffic scene consists of the autonomous vehicle (AV, controlled by PDM-Lite), background vehicles (BVs), and critical background vehicles (CBVs), which interact dynamically over time.
The visualizations demonstrate the ability of RIFT to generate temporally coherent, realistic, and controllable trajectories across a variety of traffic situations. Even under complex and evolving closed-loop conditions, RIFT maintains stable multimodal behavior, highlighting its effectiveness in simulating realistic and controllable traffic scenarios around the AV.